Tax Calculation.

Bob, a single person, is the sole owner of a business that he anticipates will have $212,000 in net taxable income per year. He is trying to decide whether to organize the business as an LLC or a C-Corporation. His attorney informed him that he’d have similar liability protection with either entity and so he is asking you about the tax consequences. Calculate the total income tax consequences for both Bob and the Business assuming all income flows through to Bob if it’s an LLC and if taxed as a corporation then all income after taxes is paid out as a dividend to Boband then draw a conclusion as to which entity has a better tax result. Assume this is Bob’s only income and that the first $12,000 of income attributable to Bob (whether as flow through income or as a dividend) is exempt from tax by Bob’s standard deduction. Use the tax calculation tables in Packet (part 1) to determine the total tax and decide which entity results in a lower income tax liability.
Answer #1

Amounts are in $

Part 1 : Incorporated as C Corporation

Firstly the income of corporation is taxed at 21% as per TCJA (Tax Cuts and Jobs Act) and next the dividend will be taxed in the hands of Individual at 15% (most dividends are taxed at 15%). Also there is 12,000 standard deduction available (amount mentioned in the question)

Tax in hands of Corporation = 21% x 212,000 = 44,520

Tax in Individual hands

Dividend = 212,000 - 44,520 = 167,480

Tax on dividend = (167,480-12,000) x 15% = 23,322

Income net of tax = 144,158

Total Income tax = 67,842

(These dividends are taken as qualified dividends so taxed at capital gain tax rates)

Part 2 : Single Person LLC

The income will have pass throuy status and is taxed directly in thr hands of Individual at applicable tax rates. Also there is 20% pass through status deduction available. The individual also have 12,000 standard deduction available.

Qualified business income = 212,000

20% pass through deduction = 42,400

Taxable Income = 169,600

Standard deduction = 12,000 (mentioned in question)

Net Taxable Income = 157,600

This is taxed at slab rates applicable

Tax = [9525 x 10% + (38700-9525) x 12% + (82500-38,700) x 22% + (157500-82500) x 24% + (157600 - 157500) x 32%]

Tax amount = 32,121.5

Income net of tax = 212,000 - 32,121.5 = 179,878.5

This shows that the income tax is lower under Single person LLC than the C Corporation thus resulting in higher income net of tax in Single person LLC.